# Novel structures and collapse of solitons in nonminimally gravitating dark matter halos

### v1

@@ Abstract
Ultralight dark matter simulations predict Bose-Einstein condensations with short-range correlation, known as solitons or boson stars, at the centers of dark matter halos. This paper investigates the formation and collapse of dark matter solitons influenced by nonminimal gravitational effects, characterized by gradient-dependent self-interactions of dark matter and an additional source in Poisson’s equation for gravity. Our simulations suggest that the initial evolution of dark matter resembles that without nonminimal gravitational effects. However, regions with negative mass density may develop, and solitons will collapse when their densities reach certain critical values for both positive and negative coupling constants. With strong nonminimal coupling, structure growth could be significantly enhanced.
@@ 1 Introduction
A popular idea of dark matter involves a sub-eV scalar field known as ultralight dark matter [1, 2, 3, 4]. This class of models can produce cosmological structures similar to cold dark matter on large scales [5, 6]. On small scales, the thermalization of scalar particles typically leads to the formation of condensations with short-range correlations called solitons or boson stars [7, 8]. Simulations in both flat and cosmological backgrounds show that solitons emerge at the centers of dark matter halos [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Solitons exhibit flat and smooth central density profiles, offering a natural solution to the discrepancy between the cuspy density profiles predicted by cold dark matter simulations and the flatter ones observed in galactic centers [2]. Ultralight dark matter is also proposed as a solution to the final parsec problem in simulations of supermassive black hole mergers [22]. In these simulations, inspirals stall at a few parsecs apart due to insufficient gravitational wave emission for orbits on those scales, which is problematic given the abundance of very massive black holes observed at galactic centers. However, orbiting near a solitonic core [23] or interacting with ultralight dark matter granules [24] could provide mechanisms to facilitate orbital decay.
@@
Solitons give rise to important opportunities for discovering and constraining ultralight dark matter. [^1] For examples, solitons might dominate the bulge dynamics in massive galaxies and could potentially represent the majority of halo mass in smaller dwarf galaxies. Analysis against observed galactic kinematics has provided strong constraints on DM mass and self-interactions [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. If dark matter is coupled to photons, solitons could generate observable electromagnetic signals [37, 38, 39, 40, 41]. A detailed understanding of their formation and growth is crucial for inferring their cosmological abundance. In the simplest scenario, where dark matter interacts only through minimal couplings to gravity, solitons can condensate from a thermal bath of dark matter particles through kinetic relaxation [42]. They continue to grow until relativistic effects become important. It is suggested that soliton growth slows once they become massive enough to heat the surrounding halo [13, 17], with the masses of the soliton $M_{s}$ and the halo $M_{h}$ approximately following the scaling $M_{s}\propto M_{h}^{1/3}$. However, there is ongoing debate about the quantitative characterization of the soliton growth rate [10, 13, 17, 43, 44]. Baryonic matter could influence this process, typically leading to more compact cores [45, 46].
@@
As dense clumps of particles, dark matter densities within solitons could be many orders of magnitude higher than average. This implies that solitons are likely to be influenced by dark matter self-interactions and the foregoing simple picture needs to be improved. These self-interactions, to the leading order in the nonrelativistic limit, are quartic in fields (cubic terms are not allowed since a single nonrelativistic particle cannot create two nonrelativistic particles and vice versa). In the case of axions—hypothetical particles proposed to explain the absence of the neutron electric dipole moment [47, 48, 49]—attractive self-interactions can destabilize solitons when their mass exceeds a threshold [50, 51, 52, 8], resulting in bursts of relativistic axions [53]. On the other hand, soliton condensation could be promoted by repulsive self-interactions [17], which can arise from spontaneous symmetry breaking and integrating out heavy degrees of freedom [54].
@@
Dark matter self-interactions can also result from their nonminimal couplings to gravity [55, 8]. The leading-order nonminimal gravitational interactions (NGIs) manifest as quartic terms in fields, including both gradient-independent and gradient-dependent parts. At first glance, the gradient-dependent part seems negligible due to the suppression by the small momenta of particles, reducing the effective self-interactions to the traditional parameterization used in previous references. However, as shown in [8], these two components are interrelated because they arise from integrating out the nonlocal gravitational potential, described by a modified Poisson’s equation. In this paper, we aim to investigate the impact of this new component of self-interactions on the formation and growth of solitons using 3+1-dimensional simulations. We will demonstrate that gradient-dependent self-interactions impose an upper bound on the mass of solitons for both positive and negative couplings, thereby affecting the mass distribution of solitons in the long-term evolution.
@@ 2 Field equations and numerical setup
where $\psi$ is the dark matter field that varies slowly in time, $m$ is the field mass, $\Phi$ is the gravitational potential, $\xi$ characterizes NGIs, and $\rho_{\xi}$ is the mass density defined as $\rho_{\xi}\equiv\rho+(\xi/m^{2})\nabla^{2}\rho$ with $\rho\equiv m|\psi|^{2}$. The natural units $c=\hbar=1$ are adopted. In this work, we focus on time scales short compared with the Hubble time, thus neglecting the expansion of the universe.
@@
To simulate the Schroedinger-Poisson equations (2.4) and (2.5), we employ a fourth-order pseudospectral method as described in [57], treating the NGI term in the Schroedinger equation as an additional part of the potential. Initial conditions are set in momentum space with a delta-like distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto\delta(|\widetilde{\bm{p}}|-v_{0}/v)$ and a random phase. Here, $v_{0}\ll 1$ is a reference velocity for the nonrelativistic system, but its exact value is unimportant due to the freedom in choosing $v$. For simplicity, we set $v=v_{0}$. Alternative initial conditions, such as a Gaussian distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto e^{-\widetilde{\bm{p}}^{2}}$ or a step-like distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto\theta(v_{0}/v-|\widetilde{\bm{p}}|)$, may also be used. However, the formation of dark matter halos and solitons is largely insensitive to the initial distribution [10]. To study isolated halos, we run simulations in a box with size $\widetilde{L}>2\pi/\widetilde{k}_{J}$, where $\widetilde{k}_{J}=(4\widetilde{\overline{\rho}})^{1/4}$ is the Jeans scale, with a bar over variables indicating their spatial mean value. The box size is set to $\widetilde{L}=18$ and the total number of particles to $\widetilde{N}=1005.3$.
@@ 3.1 Condensation of self-gravitating solitons
Figure 1: Evolution of dark matter without nonminimal gravitational effects, illustrated by snapshots of the particle number density $\widetilde{\rho}$ at different times. These snapshots show the density on a plane passing through the peak density location. We see that a dense dark matter halo forms gradually over time, with a soliton clearly emerging at the halo’s center by $\widetilde{t}\approx 110$.
@@ 3.3 Dark matter with positive nonminimal couplings
Figure 2: Evolution of dark matter with $\widetilde{\xi}=0.02$, illustrated by snapshots of the number density $\widetilde{\rho}$ (in upper panels) and the mass density $\widetilde{\rho}_{\xi}$ (in lower panels) at different times. By $\widetilde{t}=100$, a soliton has formed at the halo’s center. As the soliton grows, the mass density $\widetilde{\rho}_{\xi}$ in its inner core becomes negative when the soliton collapses at $\widetilde{t}\approx 111.3$.
@@ 3.4 Dark matter with negative nonminimal couplings
Figure 5: Evolution of dark matter with $\widetilde{\xi}=-0.008$, illustrated by snapshots of the number density $\widetilde{\rho}$ (in upper panels) and the mass density $\widetilde{\rho}_{\xi}$ (in lower panels) at different times. By $\widetilde{t}=13$, a soliton has formed at the center of the halo, surrounded by multiple subhalos. These subhalos could contain more mass than particle numbers, with regions showing negative mass density $\widetilde{\rho}_{\xi}$ in the voids between them.
@@ 4 Conclusions
We demonstrate that for both positive and negative coupling constants $\widetilde{\xi}$, solitons collapse when their densities exceed certain critical values. The mass of collapsing solitons is at the level of $\sim 40M_{\mathrm{P}}^{2}/(m|\xi|^{1/2})$ [8] . Once the critical densities are reached, the dark matter clump rapidly grows, necessitating relativistic simulations for further evolution. In contrast, solitons with solely gradient-independent self-interactions will collapse only for negative couplings, i.e., attractive self-interactions [17, 60, 61]. Collapsing solitons undergo a series of bursts and could be potentially detected if dark matter is coupled to regular matter [53].
Another novel aspect of nonminimally gravitating dark matter is the inclusion of a gradient-dependent source in Poisson’s equation, alongside the rest mass density. This leads us to consider a modified mass density $\rho_{\xi}$ that maintains the familiar Newton’s law. Our numerical results suggest that solitons with positive couplings could feature an inner core with negative mass density via dynamical matter accretion. For negative couplings, subhalos surrounding the central soliton could have more mass than particle numbers, with voids of negative mass density in between. These characteristics could have novel implications for gravitational lensing [62], dynamic heating of stars [63], and axion minivoids [64], warranting further investigation.
@@ Acknowledgments
We would like to thank Xiaolong Du and Luca Visinelli for helpful comments. JC acknowledges the support from the Fundamental Research Funds for the Central Universities and the Chongqing Natural Science Foundation General Project. HYZ is supported in part by the National Natural Science Foundation of China (NSFC) through grant No. 12350610240. This research was also facilitated by the computational resources in the School of Physical Science and Technology at Southwest University.

### v2

@@ Abstract
Ultralight dark matter simulations predict condensates with short-range correlation, known as solitons or boson stars, at the centers of dark matter halos. This paper investigates the formation and collapse of dark matter solitons influenced by nonminimal gravitational effects, characterized by gradient-dependent self-interactions of dark matter and an additional source in Poisson’s equation for gravity. Our simulations suggest that the initial evolution of dark matter resembles that without nonminimal gravitational effects. However, regions with negative potential curvature may develop, and solitons will collapse when their densities reach certain critical values for both positive and negative coupling constants. With strong nonminimal gravitational effects, we verify that linear density perturbations could grow on both large and small scales, potentially enhancing structure formation.
@@ 1 Introduction
A popular idea of dark matter involves a sub-eV scalar field known as ultralight dark matter [1, 2, 3, 4]. This class of models can produce cosmological structures similar to cold dark matter on large scales [5, 6]. On small scales, the thermalization of scalar particles typically leads to the formation of condensations with short-range correlations called solitons or boson stars [7, 8]. Focusing on the very light end of the mass spectrum $\mathcal{O}(10^{-22}$$10^{-20})\mathrm{eV}$, simulations in both expanding and nonexpanding backgrounds show that solitons emerge at the centers of dark matter halos [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. Solitons exhibit flat and smooth central density profiles, offering a natural solution to the discrepancy between the cuspy density profiles predicted by cold dark matter simulations and the flatter ones observed in galactic centers [2]. Ultralight dark matter is also proposed as a solution to the final parsec problem in simulations of supermassive black hole mergers [22]. In these simulations, inspirals stall at a few parsecs apart due to insufficient gravitational wave emission for orbits on those scales, which is problematic given the abundance of very massive black holes observed at galactic centers. However, orbiting near a solitonic core [23] or interacting with ultralight dark matter granules [24] could provide mechanisms to facilitate orbital decay. [^1] Other potential solutions to the final parsec problem involve, for example, self-interacting dark matter [25], accretion disks [26, 27, 28], or axisymmetric galactic halo profile [29] (which was debated in [30]).
Solitons give rise to important opportunities for discovering and constraining ultralight dark matter. For examples, solitons might dominate the bulge dynamics in massive galaxies and could potentially represent the majority of halo mass in smaller dwarf galaxies. Analysis against observed galactic kinematics has provided strong constraints on DM mass and self-interactions [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. If dark matter is coupled to photons, solitons could generate observable electromagnetic signals [43, 44, 45, 46, 47].
@@
Therefore, a detailed understanding about the formation and growth of solitons is crucial for inferring their cosmological abundance. In the simplest scenario, where dark matter interacts only through minimal couplings to gravity, solitons can condensate from a thermal bath of dark matter particles through kinetic relaxation [48]. They continue to grow until relativistic effects become important. It is suggested that soliton growth slows once they become massive enough to heat the surrounding halo [13, 17], with the masses of the soliton $M_{s}$ and the halo $M_{h}$ approximately following the scaling $M_{s}\propto M_{h}^{1/3}$. However, there is ongoing debate about the quantitative characterization of the soliton growth rate [10, 13, 17, 49, 50]. Baryonic matter could influence this process, typically leading to more compact cores [51, 52].
@@
As dense clumps of particles, dark matter densities within solitons could be many orders of magnitude higher than the local density $\sim 0.4\mathrm{GeV/cm^{3}}$ inferred from galactic rotation curves [53, 54]. This implies that solitons are likely to be influenced by dark matter self-interactions and the foregoing simple picture needs to be improved. These self-interactions, to the leading order in the nonrelativistic limit, are quartic in fields (cubic terms are not allowed since a single nonrelativistic particle cannot create two nonrelativistic particles and vice versa). In the case of axions—hypothetical particles proposed to explain the absence of the neutron electric dipole moment [55, 56, 57]—attractive self-interactions can destabilize solitons when their mass exceeds a threshold [58, 59, 60, 8, 61, 62, 63, 64], resulting in bursts of relativistic axions [65]. On the other hand, soliton condensation could be promoted by repulsive self-interactions [17], which can arise from spontaneous symmetry breaking and integrating out heavy degrees of freedom [66].
@@
Dark matter self-interactions can also result from their nonminimal couplings to curvature fields such as $\xi\phi^{2}R/2$, where $\phi$ is a scalar dark matter field, $R$ is the Ricci scalar, and $\xi$ characterizes the coupling strength [67, 8]. [^2] Nonminimal couplings to gravity could arise from quantum corrections and are essential for the renormalization of field theories in curved spacetime [68, 69, 70, 71, 72]. Their phenomenological implications have been studied in various contexts, such as dark matter [73, 74, 75, 76, 67], inflation [77, 78, 79, 80, 81, 82, 83, 84], and modified gravity [85, 86, 87, 88, 89, 90]. This leading-order nonminimal gravitational interaction (NGI) manifest as quartic terms in fields, including both gradient-independent and gradient-dependent parts. At first glance, the gradient-dependent part seems negligible due to the suppression by the small momenta of particles, reducing the effective self-interactions to the traditional parameterization used in previous references. However, as shown in [8], these two components are interrelated because they arise from integrating out the nonlocal gravitational potential, described by a modified Poisson’s equation. In this paper, we aim to investigate the impact of this new component of self-interactions on the formation and growth of solitons using 3+1-dimensional simulations. We will see that regions with negative mass density (defined as the source of the Poisson’s equation) could develop. By tracking the evolution of maximum densities, we demonstrate that gradient-dependent self-interactions impose an upper bound on the mass of solitons for both positive and negative couplings, thereby affecting the mass distribution of solitons in the long-term evolution.
@@ 2 Field equations and numerical setup
where $\psi$ is a nonrelativistic classical field that varies slowly in time, $m$ is the field mass, $\Phi$ is the gravitational potential, $\xi$ characterizes NGIs, and $\rho_{\xi}$ is the mass density defined as $\rho_{\xi}\equiv\rho+(\xi/m^{2})\nabla^{2}\rho$ with $\rho\equiv m|\psi|^{2}$. The natural units $c=\hbar=1$ are adopted. In this work, we focus on time scales short compared with the Hubble time, thus neglecting the expansion of the universe.
@@
To simulate the Schroedinger-Poisson equations (2.4) and (2.5), we employ a fourth-order pseudospectral method as described in [92, 93, 94], treating the NGI term in the Schroedinger equation as an additional part of the potential. Initial conditions are set in momentum space with a delta-like distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto\delta(|\widetilde{\bm{p}}|-v_{0}/v)$ and a random phase. Here, $v_{0}\ll 1$ is a reference velocity for the nonrelativistic system, but its exact value is unimportant due to the freedom in choosing $v$. For simplicity, we set $v=v_{0}$. Alternative initial conditions, such as a Gaussian distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto e^{-\widetilde{\bm{p}}^{2}}$ or a step-like distribution $|\widetilde{\psi}_{\widetilde{\bm{p}}}|^{2}\propto\theta(v_{0}/v-|\widetilde{\bm{p}}|)$, may also be used. However, the formation of dark matter halos and solitons is largely insensitive to the initial distribution [10]. To study isolated halos, we run simulations in a box with size $\widetilde{L}>2\pi/\widetilde{k}_{J}$, where $\widetilde{k}_{J}=(4\widetilde{\overline{\rho}})^{1/4}$ is the Jeans scale, with a bar over variables indicating their spatial mean value. The box size is set to $\widetilde{L}=18$ and the total number of particles to $\widetilde{N}=1005.3$.
@@ 3.1 Condensation of self-gravitating solitons
Figure 1: Evolution of dark matter without nonminimal gravitational effects, illustrated by slices of the particle number density $\widetilde{\rho}$ at different times. These slices show the density on a plane passing through the peak density location. We see that a dense dark matter halo forms gradually over time, with a soliton clearly emerging at the halo’s center by $\widetilde{t}\approx 110$.
@@ 3.3 Dark matter with positive nonminimal couplings
Figure 2: Evolution of dark matter with $\widetilde{\xi}=0.02$, illustrated by slices of the number density $\widetilde{\rho}$ (in upper panels) and the mass density $\widetilde{\rho}_{\xi}$ (in lower panels) at different times. By $\widetilde{t}=100$, a soliton has formed at the halo’s center. As the soliton grows, the mass density $\widetilde{\rho}_{\xi}$ in its inner core becomes negative when the soliton collapses at $\widetilde{t}\approx 111.3$.
@@ 3.4 Dark matter with negative nonminimal couplings
Figure 5: Evolution of dark matter with $\widetilde{\xi}=-0.008$, illustrated by slices of the number density $\widetilde{\rho}$ (in upper panels) and the mass density $\widetilde{\rho}_{\xi}$ (in lower panels) at different times. By $\widetilde{t}=13$, a soliton has formed at the center of the halo, surrounded by multiple subhalos. These subhalos could contain more mass than particle numbers, with regions showing negative mass density $\widetilde{\rho}_{\xi}$ in the voids between them.
@@ 4 Conclusions
In figures 4 and 7, we demonstrate that for both positive and negative coupling constants $\widetilde{\xi}$, solitons collapse when their densities exceed certain critical values given by equations (3.3) and (3.4). The mass of collapsing solitons is at the level of $\sim 40M_{\mathrm{P}}^{2}/(m|\xi|^{1/2})$ [8] . Once the critical densities are reached, the dark matter clump rapidly grows, necessitating relativistic simulations for further evolution. In contrast, solitons with solely gradient-independent self-interactions will collapse only for negative couplings, i.e., attractive self-interactions [17, 97, 98]. Collapsing solitons undergo a series of bursts and could be potentially detected if dark matter is coupled to regular matter [65].
Another novel aspect of nonminimally gravitating dark matter is the inclusion of a gradient-dependent source in the Poisson’s equation (2.2), alongside the rest mass density. This leads us to consider a modified mass density $\rho_{\xi}$ that maintains the familiar Newton’s law. Our numerical results suggest that solitons with positive couplings could feature an inner core with negative mass density via dynamical matter accretion, as shown in figure 2. For negative couplings, subhalos surrounding the central soliton could have more mass than particle numbers with voids of negative mass density in between, see figure 5. These characteristics could have novel implications for gravitational lensing [99], dynamic heating of stars [100], and axion minivoids [101], warranting further investigation.
@@ Acknowledgments
We would like to thank Xiaolong Du and Luca Visinelli for helpful comments. We would especially like to thank Zhipan Li for his assistance in setting up the clusters. JC acknowledges the support from the Fundamental Research Funds for the Central Universities (SWU-KR22012) and the Chongqing Natural Science Foundation General Project (2023NSCQ-MSX1929). HYZ is supported in part by the National Natural Science Foundation of China (NSFC) through grant No. 12350610240. This research was also facilitated by the computational resources in the School of Physical Science and Technology at Southwest University.